OpenStudy (anonymous):

Determine whether the sequence converges or diverges. If it converges, give the limit. (2 points) 60, 10, 5 divided by 3 , 5 divided by 18 , ...

OpenStudy (anonymous):

The sequence given is \[60, 10, \frac53, \frac5{18}, ... \]The key thing to notice here is that you're dividing by 6 each time: \[60\]\[60*\frac16 = 10\]\[10*\frac16=\frac{10}{6} = \frac53\]\[\frac53*\frac16 = \frac{5*1}{3*6} = \frac{5}{18}\] and so on. Convergence and the limit of convergence should be pretty easy to figure out from here...